# ObjectVisA-120: Object-based Visual Attention Prediction in Interactive Street-crossing Environments

### v1

@@ Abstract
The object-based nature of human visual attention is well-known in cognitive science, but has only played a minor role in computational visual attention models so far. This is mainly due to a lack of suitable datasets and evaluation metrics for object-based attention. To address these limitations, we present ObjectVisA-120  – a novel 120-participant dataset of spatial street-crossing navigation in virtual reality specifically geared to object-based attention evaluations. The uniqueness of the presented dataset lies in the ethical and safety affiliated challenges that make collecting comparable data in real-world environments highly difficult. ObjectVisA-120  not only features accurate gaze data and a complete state-space representation of objects in the virtual environment, but it also offers variable scenario complexities and rich annotations, including panoptic segmentation, depth information, and vehicle keypoints. We further propose object-based similarity (oSIM) as a novel metric to evaluate the performance of object-based visual attention models, a previously unexplored performance characteristic. Our evaluations show that explicitly optimising for object-based attention not only improves oSIM performance but also leads to an improved model performance on common metrics. In addition, we present SUMGraph, a Mamba U-Net-based model, which explicitly encodes critical scene objects (vehicles) in a graph representation, leading to further performance improvements over several state-of-the-art visual attention prediction methods. The dataset, code and models will be publicly released.
@@ II Related Works
Metrics for evaluating attention prediction. Depending on the precise task formulation, a wide variety of metrics to evaluate predictions of human visual attention are used. Methods predicting scanpaths (ordered fixation sequences), are often evaluated with sequence comparison metrics such as Dynamic Time Warping or Levenshtein similarity [18]. More relevant to our work are metrics from the field of saliency prediction, as they compare a predicted attention distribution with ground truth gaze [5, 36]. These metrics can be broadly separated into two categories. Distribution-based metrics such as Similarity (SIM), Pearson’s Correlation Coefficient (CC), and Kullback-Leibler Divergence (KLD) compare predicted with ground truth attention distributions. Location-based metrics such as Area Under the Curve (AUC) and Normalized Scanpath Saliency (NSS) compare predicted attention distributions to ground truth fixation locations. A common feature of all these metrics is that they operate on an abstract image plane without any notion of the attended objects or their semantics.
@@ III ObjectVisA-120 Dataset
TABLE I: The comparison of saliency and gaze prediction datasets in VR setups. In the column Task, we use a separate notation to represent the task, where F stands for the free viewing task, S corresponds to the search task, N stands for navigation task, SC denotes street-crossing environmental constraints, and I stands for the interaction task. The (*) symbol indicates partial or limited annotation.
@@
ObjectVisA-120 dataset. While data was recorded at 90Hz, for the purpose of this work, we chose to consider every 3rd frame, giving us a sampling rate of 30Hz, resulting in $6.142.167$ frames. In addition to the raw field-of-view RGB images (cf. Figure 1, A), the dataset comprises corresponding eye-gaze fixation information (2D/3D) and derived saliency maps (cf. Figure 1, A overlaid) computed individually, panoptic segmentation (covering vehicles, pedestrians, crossings, road and pavement surfaces, buildings, vegetation, signage, and other street furniture as unary objects), depth maps (cf. Figure 1, B-C), and skeleton labels (cf. Figure 1, D) for vehicles. Notable, since the entire state space of the study has been saved, it is possible to generate additional modalities like 2D/3D bounding boxes or add another layer of segmentation. To maintain consistency with previous work, the skeleton labelling corresponds to the sparse variation of the ApolloCar3D dataset as used in OpenPifPaf [33], which contains 24 keypoints. Incorporating pedestrian skeleton representations is beneficial due to their strong relevance for behavioural and intent inference, whereas extending graph representations to static scene elements offers limited added value and incurs unnecessary computational overhead. The panoptic segmentation on ObjectVisA-120 follows the CityScapes [10] labeling policy. The extensive labeling of ObjectVisA-120 enables multiple computer vision tasks to be performed, including human-like trajectory generation (safe versus unsafe) empowered by visual attention [51], providing a significant value to the community. Note that the annotation policy for occluded objects closely reflects real-world perception, where the occluded parts of objects are excluded from all labels (cf. Figure 1, where vehicle occluded regions are excluded from annotations, aligning with the principles of human visual perception). Where necessary, state-of-the-art methods may be used to generate on-demand segmentation labels for real-world datasets and applications. Subsequent fine-tuning on the ApolloCar3D dataset can be applied to derive the required skeleton representations of vehicles, people and animals supporting deployment in real-world settings.
To generate ground truth attention maps, we follow previous work [52], where 2D sparse ground truth attention maps were created using the last three fixation points. According to [53], the eccentric angle of the foveal region is approximately 2, additionally, the eye-tracking accuracy of HTC Vive Pro Eye  [^1] https://developer.vive.com/resources/hardware-guides/vive-pro-eye-specs-user-guide/ ranges from 0.5-$1.$ Thus, the ground-truth attention maps were generated using Gaussians with a standard deviation of $3\times dva$ (degree of visual angle).
@@ V Methodology
Loss function. Following previous work [23], our model employs a loss function that integrates several components [1, 15, 39, 23]. Additionally, we extend the loss with the introduced $oSIM$ term to optimize for object-based (safety-critical) attention explicitly:
@@ VIII Conclusion
In this work, we introduced the ObjectVisA-120 dataset, a novel VR resource applicable to diverse tasks, with a focus on visual attention prediction in street-crossing scenarios—an area challenging due to complexity, safety and ethical considerations. We proposed an object-based Similarity (oSIM) metric to capture object-driven attention, offering a perspective more aligned with human perception. Empirical results show that integrating oSIM into the loss function improves performance and advances modeling of visual attention. We also introduced the SUMGraph model, which exploits contextual graph information to improve predictive accuracy and achieve state-of-the-art performance.

### v2

@@ Abstract
The object-based nature of human visual attention is well-known in cognitive science, but has only played a minor role in computational visual attention models so far. This is mainly due to a lack of suitable datasets and evaluation metrics for object-based attention. To address these limitations, we present ObjectVisA-120  [^4] https://www.kaggle.com/datasets/igorvozniak/objectvisa-120 – a novel 120-participant dataset of spatial street-crossing navigation in virtual reality specifically geared to object-based attention evaluations. The uniqueness of the presented dataset lies in the ethical and safety affiliated challenges that make collecting comparable data in real-world environments highly difficult. ObjectVisA-120  not only features accurate gaze data and a complete state-space representation of objects in the virtual environment, but it also offers variable scenario complexities and rich annotations, including panoptic segmentation, depth information, and vehicle keypoints. We further propose object-based similarity (oSIM) as a novel metric to evaluate the performance of object-based visual attention models, a previously unexplored performance characteristic. Our evaluations show that explicitly optimising for object-based attention not only improves oSIM performance but also leads to an improved model performance on common metrics. In addition, we present SUMGraph, a Mamba U-Net-based model, which explicitly encodes critical scene objects (vehicles) in a graph representation, leading to further performance improvements over several state-of-the-art visual attention prediction methods. The dataset, code and models will be publicly released.
@@ II Related Works
Metrics for evaluating attention prediction. Depending on the precise task formulation, a wide variety of metrics to evaluate predictions of human visual attention are used. Methods predicting scanpaths (ordered fixation sequences), are often evaluated with sequence comparison metrics such as Dynamic Time Warping or Levenshtein similarity [17]. More relevant to our work are metrics from the field of saliency prediction, as they compare a predicted attention distribution with ground truth gaze [5, 34]. These metrics can be broadly separated into two categories. Distribution-based metrics such as Similarity (SIM), Pearson’s Correlation Coefficient (CC), and Kullback-Leibler Divergence (KLD) compare predicted with ground truth attention distributions. Location-based metrics such as Area Under the Curve (AUC) and Normalized Scanpath Saliency (NSS) compare predicted attention distributions to ground truth fixation locations. A common feature of all these metrics is that they operate on an abstract image plane without any notion of the attended objects or their semantics.w
@@ III ObjectVisA-120 Dataset
TABLE I: The comparison of saliency and gaze prediction datasets (VR). In the column Task, we use a separate notation to represent the task, where F stands for the free viewing task, S corresponds to the search task, N stands for navigation task, SC denotes street-crossing constraints, and I stands for the interaction task. The (*) symbol indicates partial or limited annotation.
@@
ObjectVisA-120 dataset. While data was recorded at 90Hz, for the purpose of this work, we chose to consider every 3rd frame, resulting in $6.142.167$ frames (sampling rate of 30Hz). In addition to the raw field-of-view RGB images (cf. Figure 1, A), the dataset comprises corresponding eye-gaze fixation information (2D/3D) and derived saliency maps (cf. Figure 1, A overlaid) computed individually, panoptic segmentation (covering vehicles, pedestrians, crossings, road and pavement surfaces, buildings, vegetation, signage, and other street furniture as unary objects), depth maps (cf. Figure 1, B-C), and skeleton labels (cf. Figure 1, D) for vehicles. Notable, since the entire state space of the study has been saved, it is possible to generate additional modalities like 2D/3D bounding boxes or add another layer of segmentation. To maintain consistency with previous work, the skeleton labelling corresponds to the sparse variation of the ApolloCar3D dataset as used in OpenPifPaf [32], which contains 24 keypoints. Incorporating pedestrian skeleton representations is beneficial due to their strong relevance for behavioural and intent inference, whereas extending graph representations to static scene elements offers limited added value and incurs unnecessary computational overhead. The panoptic segmentation on ObjectVisA-120 follows the CityScapes [10] labeling policy. The extensive labeling of ObjectVisA-120 enables multiple computer vision tasks to be performed, including human-like trajectory generation empowered by visual attention [52], providing a significant value to the community. Note that the annotation policy for occluded objects closely reflects real-world perception, where the occluded parts of objects are excluded from all labels (cf. Figure 1). Where necessary, state-of-the-art methods may be used to generate on-demand segmentation labels for real-world datasets and applications. Subsequent fine-tuning on the ApolloCar3D dataset can be applied to derive the required skeleton representations of vehicles, people and animals supporting deployment in real-world settings. To generate ground truth attention maps, we follow previous work [51], where 2D sparse ground truth attention maps were created using the last three fixation points. According to [53], the eccentric angle of the foveal region is approximately 2, additionally, the eye-tracking accuracy of VR headset [^5] HTC Vive Pro Eye: https://developer.vive.com/resources/hardware-guides/vive-pro-eye-specs-user-guide/ ranges from 0.5-$1.$ Thus, the ground-truth attention maps were generated using Gaussians with a standard deviation of $3\times dva$ (degree of visual angle).
@@ V Methodology
Loss function. Following previous work [22], our model employs a loss function that integrates several components [1, 15, 38, 22]. We extend the loss with the $oSIM$ term to optimize for object-based attention explicitly:
@@ VIII Conclusion
In this work, we introduced the ObjectVisA-120 dataset, a novel VR resource applicable to diverse tasks, with a focus on visual attention prediction in street-crossing scenarios, an area challenging due to complexity, safety and ethical considerations. We proposed an object-based Similarity (oSIM) metric to capture object-driven attention, offering a perspective more aligned with human perception. Empirical results show that integrating oSIM into the loss function improves performance and advances modeling of visual attention. We also introduced the SUMGraph model, which exploits contextual graph information to improve predictive accuracy and achieve state-of-the-art performance.
IX Acknowledgment
This work was supported by the European Regional Development Fund (EFRE; EFRE-AuF-0000866) and in part by the European Union’s Horizon Europe research and innovation programme (No. 101076360). The initial study recording has been funded by the German Ministry for Research and Education (BMBF; 01IW17003), and partly by the New Energy and Industrial Technology Development (NEDO; JPNP18010).